What is the molar concentration of Fe2+ ion in an aqueous solution if 35.50 mL of 0.113 M KBrO3 is required for complete reaction with 13.00 mL of the Fe2+ solution? The net ionic equation is:
6Fe2+(aq)+BrO3−(aq)+6H+(aq)→6Fe3+(aq)+Br−(aq)+3H2O(l).
1
Expert's answer
2020-10-05T14:26:14-0400

Solution:

For redox reaction

M1 × n1 × V1 = M2 × n2 × V

M - molarity

M Fe2+ × n Fe2+ × V Fe2+ = M KBrO3 × n KBrO3 × V KBrO3

For the following reaction calculate molarity M Fe2+

n of KBrO3 = 1; n of Fe 2+ = 1/6

M Fe 2+ = M KBrO3 × n KBrO3 × V KBrO3 / n Fe2+ × V Fe2+ = 0.113×1×35.50 / 1/6×13.0 = 1.85M


Answer: Molarity of the Fe2+ solution is 1.85M.